# Zebrafix: Mitigating Memory-Centric Side-Channel Leakage via Interleaving

### v2

@@ VII-B Comparison of Silent Store Mitigations
The Zebrafix implementation protects against leakage from silent stores without requiring further adaptions as long as the silencing of writes cannot be decomposed into smaller chunks. We therefore compare Zebrafix to the silent store mitigation cio [15] that has also been implemented for x86-64. The Zebrafix implementation addresses the general root cause of memory-centric leakages. In line with that, the proof-of-concept implementation does not include a tracking for leaking or secret-dependent instructions, but rather instruments all memory writes. The silent store-only version of cio only covers a subset of all writes that are instrumented in Zebrafix. Zebrafix does, in contrast to cio, not have an automated verification mechanism but rather introduces manual verification of the non-leakiness of memory writes that are introduced by the register allocation.
@@
The performance results for cio are based on results reported in their paper. The Zebrafix measurement setup equals the one described in Section VI-B. In line with cio, we use the $\mathtt{libsodium}$ reference implementations and protect stores of up to 8 byte of data. Protecting a binary against silent store leakage with cio introduces an overhead of $1.85$$\times$ for $\mathtt{ChaCha20}$ and an overhead of $3.76$$\times$ for $\mathtt{Ed25519}$. In contrast to that, the ciphertext side-channel protection from Zebrafix introduces an overhead factor of $2.1$$\times$ for $\mathtt{ChaCha20}$ and a factor of $1.6$$\times$ for $\mathtt{Ed25519}$.
@@
The memory tracking Pintool without further adaption gives an overapproximation of the leakages through silent stores because if we can rule out leakages through any collision, we can also rule out leakages through consecutive memory writes leading to a collision. The overapproximation can be refined if the Pintool is often needed for verifying implementations in the silent store scenario.
@@ X Conclusion
In this work, we have studied memory-centric side-channel leakages that render constant-time code guarantees ineffective. We have discussed prerequisites for efficient ciphertext side-channel mitigations and implemented Zebrafix. We have shown that Zebrafix and interleaving in general can improve both the performance and security guarantees of ciphertext side-channel mitigations, though this may come with practical constraints such as source code adjustments. Moreover, we have leveraged interleaving to devise mitigations against more generic memory-centric side-channel leakages like silent stores, and explained how holistic defenses can be developed. We conclude that, despite its limitations, interleaving is a promising approach to counter ciphertext side-channels and silent stores.
@@ Acknowledgment
This work has been supported by Deutsche Forschungsgemeinschaft (DFG) through the ReTEE project, and by Bundesministerium für Bildung und Forschung (BMBF) through the ENCOPIA project.

### v3

@@ VII-B Comparison of Silent Store Mitigations
The Zebrafix implementation protects against leakage from silent stores without requiring further adaptions as long as the silencing of writes cannot be decomposed into smaller chunks. We therefore compare Zebrafix to the silent store mitigation cio [15] that has also been implemented for x86-64. The Zebrafix implementation addresses the general root cause of memory-centric leakages. In line with that, the proof-of-concept implementation does not include a tracking for leaking or secret-dependent instructions, but rather instruments all memory writes. The silent store-only version of cio only covers a subset of all writes that are instrumented in Zebrafix. Zebrafix does, in contrast to cio, not have an automated verification mechanism but rather introduces manual assessment of the non-leakiness of memory writes that are introduced by the register allocation.
@@
The performance results for cio are based on results reported in their paper. The Zebrafix measurement setup equals the one described in Section VI-B. In line with cio, we use the $\mathtt{libsodium}$ reference implementations and protect stores of up to 8 byte of data. Protecting a binary against silent store leakage with cio introduces an overhead of 1.851.851.85$\times$ for $\mathtt{ChaCha20}$ and an overhead of 3.763.763.76$\times$ for $\mathtt{Ed25519}$. In contrast to that, the ciphertext side-channel protection from Zebrafix introduces an overhead factor of 2.12.12.1$\times$ for $\mathtt{ChaCha20}$ and a factor of 1.61.61.6$\times$ for $\mathtt{Ed25519}$.
@@
The memory tracking Pintool without further adaption gives an overapproximation of the leakages through silent stores because if we can rule out leakages through any collision, we can also rule out leakages through consecutive memory writes leading to a collision. The overapproximation can be refined if the Pintool is often needed for assessing implementations in the silent store scenario.
@@ X Conclusion
In this work, we have studied memory-centric side-channel leakages that render constant-time code guarantees ineffective. We have discussed prerequisites for efficient ciphertext side-channel mitigations and implemented Zebrafix. We have shown that interleaving can improve both the performance and security guarantees of ciphertext side-channel mitigations, though this may come with practical constraints such as source code adjustments. Moreover, we have leveraged interleaving to devise mitigations against more generic memory-centric side-channel leakages like silent stores, and explained how holistic defenses can be developed. We conclude that, despite its limitations, interleaving is a promising approach to counter ciphertext side-channels and silent stores.
@@ Acknowledgment
We would like to thank the anonymous reviewers for their thorough assessment and providing helpful suggestions for improvement. We also thank Johannes Liebenow and Pajam Pauls for proofreading and valuable feedback. This work has been supported by Deutsche Forschungsgemeinschaft (DFG) through the ReTEE project, and by Bundesministerium für Bildung und Forschung (BMBF) through the ENCOPIA project.